OpenStudy (anonymous):

Find the third degreeTaylor polynomial centered at 1 for the function f(x)=x+3/x

OpenStudy (anonymous):

Recall the formula for the Taylor Series/Maclaurin Series for \(f(x)\) centered at \(x=c\). Here, \(f^{(n)}(x)\) denotes the \(n\)th derivative of \(f\). Keep in mind that this equality is only valid within the radius of convergence of the series. Also, note that \(0!=1\).\[f(x)=\sum_{n=0}^\infty \frac{f^{(n)(c)}}{n!}(x-c)^n\]The third degree series is the one containing powers of \(x\) up to 3. Thus, plug and chug the following for \(f(x)=\frac{x+3}{x}\). Holler if you need more help!\[f(x) \approx f(c)+f'(x)(x-c)+\frac{f''(c)}{2}(x-c)^2+\frac{f'''(c)}{6}(x-c)^3\]Keep in mind that this polynomial is only an approximation as it contains 4 terms instead of an infinite number of terms.

OpenStudy (anonymous):

so what value do I put in for c